The ratio of Zoya’s age to Nelya’s age is 1:4. Twenty years from now, Nelya will be twice as old as Zoya will be then. Find their present ages.
Answer by ankor@dixie-net.com(22740)

The ratio of Zoya’s age to Nelya’s age is 1:4.
let x = the multiplier
then
x = Z's age
and
4x = N's age
:
Twenty years from now, Nelya will be twice as old as Zoya will be then.
4x + 20 = 2(x+20)
4x + 20 = 2x + 40
4x - 2x = 40 - 20
2x = 20
x = 20/2
x= 10 is the multiplier, there
Z is 10 and N is 40
:
:
Check solution in the statement
"Twenty years from now, Nelya will be twice as old as Zoya will be then."
40 + 20 = 2(10 + 20)
60 = 2(30)